# Discrete homotopy and homology theories for finite posets

### v1

@@ 3. Discrete homotopy theory for fintie posets
    $(w,q_{1}-1)\in\mathbb{Z}_{\leq r_{1}}^{n}\times(q_{1}-1)$, $(w^{\prime},q_{1})\in\mathbb{Z}_{\leq r_{1}}^{n}\times q_{1}$ and $(w^{\prime},q_{1})\leq(w,q_{1}-1)$.
  • (3)
    $(w,0)\in\mathbb{Z}_{\leq r_{1}}^{n}\times 0$, $(w^{\prime},1)\in\mathbb{Z}_{\leq r_{1}}^{n}\times 1$ and $(w,0)\leq(w^{\prime},1)$.
For case (1), $(\lambda_{r_{1}}(w),t)\leq(\lambda_{r_{1}}(w^{\prime}),t^{\prime})$ indicates that
$\lambda_{r_{1}}(|\mathcal{K}(U^{\gamma_{r_{1}}}_{w})|)\times\beta_{q_{1}}(|\mathcal{K}(U^{I_{q_{1}}}_{t})|)\subseteq\lambda_{r_{1}}(|\mathcal{K}(U^{\gamma_{r_{1}}}_{w^{\prime}})|)\times\beta_{q_{1}}(|\mathcal{K}(U^{I_{q_{1}}}_{t^{\prime}})|),$
then by the definition of $H$, $H(\lambda_{r_{1}}(w),t)\leq H(\lambda_{r_{1}}(w^{\prime}),t^{\prime})$. For case (2), since $x_{0}$ is a minimum vertex, $x_{0}\leq H(w^{\prime},q_{1})\leq H(w,q_{1}-1)$. For case (3), we have
$H(w,0)$ $=\widehat{\mu_{X}}\circ(|\mathcal{K}(f)|\circ\lambda_{r}^{-1})\circ\lambda_{r_{1}}(w)=\widehat{\mu_{X}}\circ G\circ(\lambda_{r_{1}}(w),0)$
$\subseteq\widehat{\mu_{X}}\circ G\circ\lambda_{r_{1}}(|\mathcal{K}(U^{\gamma_{r_{1}}}_{w^{\prime}})|)\times\beta_{q_{1}}(|\mathcal{K}(U^{I_{q_{1}}}_{1})|)\subseteq U_{u}^{X},$
where $H(w^{\prime},1)=u$. This says that $H(w,0)\leq H(w^{\prime},1)$.
We claim that
(3.4) $[H_{0}]=[f]$
in $\pi_{n}^{D}(X,x_{0})$. There is a sequence of partitions
$P_{r_{1}}=A_{k+1}\subseteq A_{k}\subseteq\cdots\subseteq A_{1}\subseteq A_{0}=P_{r+1},$
the former one being a refinement of the latter one by adding an $(n-1)$-dimensional hyperplane. For each $i$, construct an order-preserving map $F_{i}:(\mathbb{Z}^{n},\partial\mathbb{Z}^{n})\rightarrow(X,x_{0})$ by
$F_{i}(w)=\begin{cases}\widehat{\mu_{X}}\circ(|\mathcal{K}(f)|\circ\lambda_{r}^{-1})\circ\lambda_{A_{i}}(w)~{}&\text{if}~{}w\in\gamma_{A_{i}},\\ x_{0}~{}&\text{otherwise},\end{cases}$
where $11\leq i\leq k+$ 1. If we can show that
(3.5) $[F_{i}]=[F_{i-1}],$
@@
$[H_{0}]=[F_{k+1}]=\cdots=[F_{0}]=[f],$
@@
which the claim follows. Without loss of generality we may assume the $(n-1)$-dimensional hyperplane added to $A_{i}$ from $A_{i-1}$ is $x_{1}=\frac{1}{r_{1}}$. Let $F^{j}:(\mathbb{Z}^{n},\partial\mathbb{Z}^{n})\rightarrow(X,x_{0})$ be defined by
$(x_{1},x_{2},\cdots,x_{n})\mapsto\begin{cases}F_{i-1}(j,x_{2},\cdots,x_{n})~{}&\text{if}~{}x_{1}=j~{}\text{or}~{}j+1,\\ F_{i-1}(x_{1},x_{2},\cdots,x_{n})~{}&\text{if}~{}x_{1}<j,\\ F_{i-1}(x_{1}-1,x_{2},\cdots,x_{n})~{}&\text{if}~{}x_{1}>j+1,\end{cases}$
where $11\leq j\leq r+$ 1. Then
$F_{i}=F^{1}>F^{2}<F^{3}>\cdots>(<)F^{r+1}=F_{i-1}.$
Thus $[f]=[H_{0}]=0$ in $\pi_{n}^{D}(X,x_{0})$, finishing the proof of the injectivity of $\theta$.
Let $f\in\pi_{n}(|\mathcal{K}(X)|,x_{0})$. The key observation now is that under the homeomorphism $\lambda_{r}$, $\mathcal{K}(\gamma_{r})$ becomes a triangulation of $L^{n}$. To check surjectivity of $\theta$, represent $f$ by a simplicial map $f^{\prime}$ from $\mathcal{K}(\gamma_{r})$ to $\mathcal{K}(X)$ for some $r>0$, using the proof of simplicial approximation theorem [10, Lemma 14.1, Theorem 16.1]. As above, we can get an order-preserving map $g:(\mathbb{Z}^{n},\partial\mathbb{Z}^{n})\rightarrow(X,x_{0})$ given by
@@
$g(w)=\begin{cases}\widehat{\mu_{X}}\circ(|f^{\prime}|\circ\lambda_{r}^{-1})\circ\lambda_{r_{1}}(w)~{}&\text{if}~{}w\in\mathbb{Z}_{\leq r_{1}}^{n},\\ x_{0}~{}&\text{otherwise},\end{cases}$
with $r_{1}$ a multiple of $r$. $\theta(g)$ is homotopic to $|\mathcal{K}(f^{\prime})|$, because both of them agree on the boundary of each cube of $\lambda_{r}(|\mathcal{K}(\gamma_{r})|)$. ∎
@@
We now directly compute the $1$th discrete homotopy group of a circle in the next example, leading to a new viewpoint toward the fundamental group of a circle.
@@
${c}$${d}$${a\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${b\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$
@@
where $x_{i}$ is the value of $f$ at $i$. If two consecutive points $x_{i}$ and $x_{i+1}$ are equal, $-r+1\leq i\leq r-2$, for example, $x_{i}=x_{i+1}$ for $20\leq i\leq r-$ 2 and $i$ is even, then $f\geq f_{1}$, where $f_{1}\in\pi^{D}_{1}(X,b)$ is defined by
@@
where $m,n<r$, $y_{-n+1},~{}y_{m-1}\not=b$, and two consecutive points $y_{i}$ and $y_{i+1}$, $-n\leq i\leq m-2$, are not equal.
@@
which means $[f]=0\in\pi^{D}_{1}(X,b)$. We now assume the values of $f$, thus $g$, contain $a$. If $y_{i}=y_{i+2}=a$, then
@@
where if $z_{i}=a$ for $-n+2\leq i\leq m-2$, then $z_{i-2}=z_{i+2}=b$, and if $z_{i}=b$ for $-n+4\leq i\leq m-4$, then $z_{i-2}=z_{i+2}$, and two consecutive vertices $z_{i}$ and $z_{i+1}$, $-n+1\leq i\leq m-2$, are distinct.
@@
where $-n+1\leq i\leq m-3$. If $z_{i}$ and $z_{i-1}$ are equal, for example, $z_{i}=z_{i+2}=c$, then
@@
$[f]=[k]=(t_{2}-t_{1})[e].$
@@ 4.1. Discrete cubical homology
Given a finite poset $M$. Let $Q_{1}$ denote the finite poset with vertex set $\{0,1\}$ and an edge $(0,1)$. For each $n\geq 1$, let $Q_{n}$ be the finite poset given by
@@
An $n$-cube of $M$ is a map $f:Q_{n}\rightarrow M$.
Let $L_{n}(M)$ be the free abelian group generated by all $n$-cubes. We define two face maps from $L_{n}(M)$ to $L_{n-1}(M)$. Let $\sigma\in L_{n}(M)$. For $n\geq 1$, $i\in\{1,\cdots,n\}$ and $(a_{1},\cdots,a_{n-1})\in Q_{n-1}$, let $f_{i}^{-}\sigma$ and $f_{i}^{+}\sigma$ be the $(n-1)$-cubes of $M$ given by
@@
An $n$-cube $\sigma$ is degenerate if $f_{i}^{-}=f_{i}^{+}$ for some $i$. Let $D_{n}(M)$ be the free abelian group generated by all degenerate $n$-cubes and $C_{n}(M)=L_{n}(M)/D_{n}(M)$, whose elements are called $n$-chains in $M$.
@@
The boundary of an $n$-cube $\sigma:Q_{n}\rightarrow M$ is
@@
The boundary operator extends to a group homomorphism $\partial_{n}^{\rm Cube}:L_{n}(M)\rightarrow L_{n-1}(M)$. Since $\partial_{n}^{\rm Cube}(D_{n}(M))\subseteq D_{n-1}(M)$, we obtain a map $\partial_{n}^{\rm Cube}:C_{n}(M)\rightarrow C_{n-1}(M)$ and a chain complex $(C_{*}(M),\partial_{*}^{\rm Cube})$. Let
$H^{\rm Cube}_{n}(M)=\rm{Ker}\partial_{n}^{\rm Cube}/\rm{Im}\partial_{n+1}^{\rm Cube}.$
The group $H_{n}^{\rm Cube}(M)$ is called the $n$th discrete cubical homology group of $M$. If $z\in\rm{Ker}\partial_{n}^{\rm Cube}$, then we write $[z]$ for the equivalence class of $z$ in $H_{n}^{\rm Cube}(M)$.
@@
Let $M_{1}$ and $M_{2}$ be finite posets. If $M_{1}$ and $M_{2}$ are homotopy equivalent, then
$H^{\rm Cube}_{n}(M_{1})=H^{\rm Cube}_{n}(M_{2})$
@@
The idea of the proof of this theorem is exacly the same as the one in [3, Theorem 3.4]. We will not prove this theorem here.
@@
Let $K$ be a simplicial complex. For $n\geq 0$, we denote by $L^{\rm Simpl}_{n}(K)$ the free abelian group generated by all nondegenerate simplices of $K$ and denote by $D^{\rm Simpl}_{n}(K)$ the free abelian group generated by all degenerate simplices of $K$. Let $C^{\rm Simpl}_{n}(K)=L^{\rm Simpl}_{n}(K)/D^{\rm Simpl}_{n}(K)$. The $n$th simplicial homology group of $K$ is denoted $H_{n}^{\rm Simpl}(K)$ whose associated chain complex is $(C^{\rm Simpl}_{*}(K),\partial_{*}^{\rm Simpl})$. Similarly, $H_{*}^{\rm Sing}$ denotes the singular homology theory. Denote by $\eta_{*}$ the canonical homomorphism
@@
To establish a natural homomorphism between the homology groups ${H}_{*}^{\rm Cube}(M)$ and ${H}_{*}^{\rm Simpl}(\mathcal{K}(M))$, we are motivated by the construction of the one in [3], which is a natural map from the cubical homology groups to the path homology groups. This map will be defined in a similar fashion.
Consider an $n$-cube $\sigma:Q_{n}\rightarrow M$ with $n\geq 1$. In order to define a map from ${C}_{n}^{\rm Cube}(M)$ to ${C}^{\rm Simpl}_{n}(\mathcal{K}(M))$, we associate to any permutation $\tau\in S_{n}$ an $n$-chain $p_{\tau}=\{x_{0},x_{1},\cdots,x_{n}~{}|~{}x_{0}<x_{1}<\cdots<x_{n}\}$ from $(0,\cdots,0)\in V(Q_{n})$ to $(1,\cdots,1)\in V(Q_{n})$. The chain $p_{\tau}$ is defined as follows. Let $x_{0}=(0,\cdots,0)$. For $0<i\leq n$, flipping the $\tau(i)$-th coordinate of $x_{i-1}$ from 00 to $1$ gives $x_{i}$. We write $p_{\tau}(i)$ for the $i$-th vertex in the chain $p_{\tau}$ for $0\leq i\leq n$.
To each $n$-cube $\sigma\in{L}_{n}^{\rm Cube}(M)$ for $n\geq 1$, we assign the chain
@@
of ${L}^{\rm Simpl}_{n}(\mathcal{K}(M))$. It is not hard to see that if $\sigma\in{D}_{n}^{\rm Cube}(M)$, then $\psi(\sigma)\in{D}^{\rm Simpl}_{n}(\mathcal{K}(M))$.
The proof that $\psi$ is a chain map is similar to the one in [3, Lemma 5.1]. For completeness, we will sketch this proof.
@@
Let $\sigma\in{C}_{n}^{\rm Cube}(M)$. Then
@@
$\psi_{*}:{H}_{p}^{\rm Cube}(M)\rightarrow{H}_{p}^{\rm Simpl}(\mathcal{K}(M))$
@@
Let $11\leq l\leq n-$ 1 . If $\tau^{\prime}$ is constructed from $\tau$ by interchanging $\tau(l)$ and $\tau(l+1)$, then $p_{\tau}(i)=p_{\tau^{\prime}}(i)$ for $i\not=l$. Thus the $l$-th summands of (4.1) for $p_{\tau}$ and $p_{\tau^{\prime}}$ coincide. In addition, ${\rm sign}(\tau)={\rm-sign}(\tau^{\prime})$. This means that
@@
Let $M$ be a homogeneous poset of dimension $1$. The map $\psi$ induces an isomorphism $\psi_{*}:{H}_{p}^{\rm Cube}(M)\rightarrow{H}_{p}^{\rm Simpl}(\mathcal{K}(M))$ for any $p\geq 0$.

### v2

@@ 3. Discrete homotopy theory for finite posets
    $|\sigma|_{\epsilon}$ contains no other vertices of $\mathcal{K}(X)$ besides those of $\sigma$.
Cover $L^{n}$ by the open sets
$|f^{\prime}|^{-1}(|\sigma|_{\epsilon})$
as $\sigma$ ranges over all maximal simplices of $\mathcal{K}(X)$. Denote this open cover by $\mathcal{A}$. Let $\delta_{\mathcal{A}}$ be the Lebesgue number for this open cover $\mathcal{A}$. We take $r_{1}$ to be an even multiple of $r$ so that each subpolyhedron of
$\lambda_{r_{1}}(|\mathcal{K}(\gamma_{r_{1}})|)$
has diameter less than $\frac{\delta}{2}$ where $\delta<\delta_{\mathcal{A}}$. By the choice of $r_{1}$, $\mathcal{K}(\gamma_{r_{1}})$ can be regarded as a refinement of $\mathcal{K}(\gamma_{r})$.
Consider a map $g:(\mathbb{Z}^{n},\partial\mathbb{Z}^{n})\rightarrow(X,x_{0})$ given by
$g(w)=\begin{cases}{\widetilde{\mu}_{X}}\circ J_{T_{r_{1}}}(|f^{\prime}|)(w),~\text{if}~w\in\gamma_{{r_{1}}},\\ x_{0},~\text{otherwise.}\end{cases}$
$g$ is order-preserving by Lemma 3.4 and Lemma 3.5. We claim that
(3.5) $\mu_{X}\circ|\mathcal{K}(g)|\circ\lambda^{-1}_{r_{g}}]=[\mu_{X}\circ|\mathcal{K}(g)|\circ\lambda^{-1}_{r_{1}}]$
in $\pi_{n}(X,x_{0})$, the classical $n$-th homotopy group of $X$. For an arbitrary partition $T$ of $L^{n}$, there is a sequence of partitions
(3.6) $T=A_{d}\supseteq\cdots\supseteq A_{k}\supseteq A_{k-1}\supseteq\cdots\supseteq A_{1}\supseteq A_{0}=T_{1},$
the former one being a refinement of the latter one by adding an (n1)(n-1)-dimensional hyperplane. For each $k$, construct a continuous map
$F_{k}:(L^{n},\partial L^{n})\rightarrow(X,x_{0})$
by
$F_{k}(w)=\mu_{X}\circ|\mathcal{K}(g)|\circ\lambda^{-1}_{A_{k}}(w)$
where $0\leq i\leq d$. If we can show that
(3.7) $F_{k}]=[F_{k-1}],$
@@
(3.8) $F_{d}]=\cdots=[F_{0}]=[\mu_{X}\circ|\mathcal{K}(g)|\circ\lambda^{-1}_{1}].$
Denote by
$1=a_{-m^{\prime}}<a_{-m^{\prime}+1}<\cdots<a_{0}<\cdots<a_{m}=1$
@@
the $1$-st component of the partition $A_{k-1}$. Without loss of generality we may assume the (n1)(n-1)-dimensional hyperplane added to $A_{k}$ from $A_{k-1}$ is $x_{1}=\tilde{a}_{0}$, where $a_{0}<\tilde{a}_{0}<a_{1}$. Let
$S:(L^{n},\partial L^{n})=\left(\lambda_{A_{k}}(|\mathcal{K}(\gamma_{A_{k}})|),\partial\lambda_{A_{k}}(|\mathcal{K}(\gamma_{A_{k}})|)\right)\rightarrow(X,x_{0})$
be the continuous map defined by
$(x_{1},x_{2},\cdots,x_{n})\mapsto\begin{cases}F_{k}(a_{0},x_{2},\cdots,x_{n}),~\text{if}~x_{1}\in[a_{0},\tilde{a}_{0}],\\ F_{k}(x_{1},x_{2},\cdots,x_{n}),~\text{if}~x_{1}<a_{0},\\ F_{k}\lvert[a_{0},\tilde{a}_{0}],~\text{if}~x_{1}\in[\tilde{a}_{0},a_{1}],\\ F_{k}\lvert[\tilde{a}_{0},a_{1}],~\text{if}~x_{1}\in[a_{1},a_{2}],\\ F_{k}\lvert[a_{i-1},a_{i}],~\text{if}~x_{1}\in[a_{i},a_{i+1}],~i>1.\end{cases}$
Checking through the definitions, one see that $F_{k-1}$ is equal to $S$ at all vertices of $\lambda_{A_{k}}(|\mathcal{K}(\gamma_{A_{k}})|)$ except at the vertex
$(\tilde{a}_{0},x_{2},\cdots,x_{n}).$
Then by the choice of $r_{1}$, we have that the following two points
$b_{1}=(\lambda_{r_{1}})\circ\lambda^{-1}_{A_{k-1}}(\tilde{a}_{0},x_{2},\cdots,x_{n}),$
$b_{2}=(\lambda_{r_{1}})\circ\lambda^{-1}_{A_{k}}(a_{0},x_{2},\cdots,x_{n}),$
lie in $|f^{\prime}|^{-1}(|\sigma|_{\epsilon})$ for some maximal simplex $\sigma$ of $\mathcal{K}(X)$. Note that $f^{\prime}$ takes simplices to simplices. With this in mind one can draw a stronger conclusion that $|f^{\prime}|(b_{1})$ and $|f^{\prime}|(b_{2})$ are indeed contained in $|\sigma|$, since $b_{1}$ and $b_{2}$ lie in some simplex of $K$, $\mathcal{K}(\gamma_{r_{1}})$ being regarded as a refinement of $K$. This means that $\mu_{X}\circ|f^{\prime}|(b_{1})$ and $\mu_{X}\circ|f^{\prime}|(b_{2})$ are comparable, all vertices of a simplex being pairwise comparable, which is equivalent to saying that $F_{k-1}$ and $S$ are comparable at the point $(\tilde{a}_{0},x_{2},\cdots,x_{n}).$ Hence we can learn that
(3.9) $F_{k-1}]=[S]$
Having taken care of the comparability between $F_{k-1}$ and $S$, we turn now to show that $F_{k}$ and $S$ are also comparable. Let
$S_{0}:(L^{n},\partial L^{n})=\left(\lambda_{A_{k}}(|\mathcal{K}(\gamma_{A_{k}})|),\partial\lambda_{A_{k}}(|\mathcal{K}(\gamma_{A_{k}})|)\right)\rightarrow(X,x_{0})$
be the continuous map defined by
$(x_{1},x_{2},\cdots,x_{n})\mapsto\begin{cases}F_{k}(\tilde{a}_{0},x_{2},\cdots,x_{n}),~\text{if}~x_{1}\in[\tilde{a}_{0},a_{1}],\\ F_{k}(x_{1},x_{2},\cdots,x_{n}),~\text{if}~x_{1}<\tilde{a}_{0},\\ F_{k}\lvert[\tilde{a}_{0},a_{1}],~\text{if}~x_{1}\in[a_{1},a_{2}],\\ F_{k}\lvert[a_{i-1},a_{i}],~\text{if}~x_{1}\in[a_{i},a_{i+1}],~i>1.\end{cases}$
Similarly, when $i>0$, let
$S_{i}:(L^{n},\partial L^{n})=\left(\lambda_{A_{k}}(|\mathcal{K}(\gamma_{A_{k}})|),\partial\lambda_{A_{k}}(|\mathcal{K}(\gamma_{A_{k}})|)\right)\rightarrow(X,x_{0})$
be the continuous map given by
$(x_{1},x_{2},\cdots,x_{n})\mapsto\begin{cases}F_{k}(a_{i},x_{2},\cdots,x_{n}),~\text{if}~x_{1}\in[a_{i},a_{i+1}],\\ F_{k}(x_{1},x_{2},\cdots,x_{n}),~\text{if}~x_{1}<a_{i},\\ F_{k}\lvert[a_{i},a_{i+1}],~\text{if}~x_{1}\in[a_{i+1},a_{i+2}].\end{cases}$
Denote $S_{-1}=S$. When $1\leq i\leq m-1$, consider the relation between $S_{i}$ and $S_{i+1}$. The proof that $F_{k-1}$ and $S_{-1}$ is comparable still works in this context, so
(3.10) $S]=[S_{-1}]=[S_{0}]=\cdots=[S_{m}]=[F_{k}].$
In view of (3.9) and (3.10), the proof of (3.7) is finished.
Now let $T=r_{1}$ and $T=T_{r_{g}}$ in (3.6) respectively. Then by (3.8)
$\mu_{X}\circ|\mathcal{K}(g)|\circ\lambda^{-1}_{r_{1}}]=[\mu_{X}\circ|\mathcal{K}(g)|\circ\lambda^{-1}_{1}]=[\mu_{X}\circ|\mathcal{K}(g)|\circ\lambda^{-1}_{r_{g}}],$
which finishes the proof of the claim (3.5). Note that
$\mu_{X}\circ|f^{\prime}|]=[\mu_{X}\circ|\mathcal{K}(g)|\circ\lambda^{-1}_{r_{1}}].$
Since $\mu_{X}$ is a weak homotopy equivalence [1, Theorem 1.4.6], the equality (3.5) leads us to conclude that
$\theta(g)=[|f^{\prime}|]=[f].$
The proof of the surjectivity of $\theta$ is complete.
For injectivity of $\theta$, suppose $h:(\mathbb{Z}^{n},\partial\mathbb{Z}^{n})\rightarrow(X,x_{0})$ is a map of radius $r_{h}$ such that $\theta([h])=0$. Let $G:(L^{n},\partial L^{n})\times I\rightarrow(|\mathcal{K}(X)|,x_{0})$ is a homotopy from $|\mathcal{K}(h)|\circ\lambda_{r_{h}}^{-1}$ to the constant map $C_{x_{0}}$. Since $L^{n}\times I$ can be viewed as a subspace of $L^{n+1}$, the preceding argument can be applied to the map $G$ with $G$ in place of $f$ in the current situation. Thus we can obtain a simplicial map $G^{\prime}$ and a homotopy $H$ corresponding to $f^{\prime}$ and $g$ respectively.
The homotopy $H$ takes the form
$H:(\mathbb{Z}^{n},\partial\mathbb{Z}^{n})\times I_{p}\rightarrow(X,x_{0}),$
@@
$(w,t)\mapsto\begin{cases}{\widetilde{\mu}_{X}}\circ J_{T_{r_{1}}}(|G^{\prime}|)(w,t),~\text{if}~(w,t)\in\gamma_{{r_{1}}}\times I_{p},\\ x_{0},~\text{otherwise,}\end{cases}$
where $p\geq 1$. Notice that $|\mathcal{K}(h)|\circ\lambda_{r_{h}}^{-1}$ is already simplicial on the underlying simplicial complex of $\lambda_{r_{h}}(|\mathcal{K}(\gamma_{r_{h}})|)$. Then by the proof in simplicial approximation theorem [15, Theorem 16.1], $G^{\prime}$ can be chosen so that the relationship between $H_{0}$ and $h$ is essentially the same as the relationship between $F_{k}$ and $S$ in the proof for the surjectivity part. Thus the similar argument shows that
$H_{0}]=[h].$
Then $[h]=0$ in $\pi_{n}^{D}(X,x_{0})$, proving the injectivity of $\theta$. ∎
To get a feeling for how the discrete homotopy groups can be derived combinatorially, let us first look at a simple example.
Example 3.7.
If $X$ is a poset with a maximum or minimum $x_{0}$, then all discrete homotopy groups vanish. To see this, note that any map $f:\mathbb{Z}^{n}\rightarrow X$ is comparable with the constant map $C_{x_{0}}:\mathbb{Z}^{n}\rightarrow x_{0}$. More precisely, either $f\leq C_{x_{0}}$ or $f\geq C_{x_{0}}$ occurs. For example, the poset $X$ depicted below has a maximum $d$ and therefore $\pi_{n}^{D}(X,d)=0$ for all $n$.
${d}$${b\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${c\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${a\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$.
@@
We now directly compute the $1$-st discrete homotopy group of a circle in the next example, leading to a new viewpoint towards the fundamental group of a circle.
@@
[…]
@@
where $x_{i}$ is the value of $f$ at $i$. If two consecutive points $x_{i}$ and $x_{i+1}$ are equal, $r+1\leq i\leq r-2$, for example, $x_{i}=x_{i+1}$ for $0\leq i\leq r-2$ and $i$ is even, then $f\geq f_{1}$, where $f_{1}\in\pi^{D}_{1}(X,b)$ is defined by
@@
where $m,n<r$, $y_{-n+1},~y_{m-1}\not=b$, and two consecutive points $y_{i}$ and $y_{i+1}$, $n\leq i\leq m-2$, are not equal.
@@
which means $f]=0\in\pi^{D}_{1}(X,b)$. We now assume the values of $f$, thus $g$, contain $a$. If $y_{i}=y_{i+2}=a$, then
@@
where if $z_{i}=a$ for $n+2\leq i\leq m-2$, then $z_{i-2}=z_{i+2}=b$, and if $z_{i}=b$ for $n+4\leq i\leq m-4$, then $z_{i-2}=z_{i+2}$, and two consecutive vertices $z_{i}$ and $z_{i+1}$, $n+1\leq i\leq m-2$, are distinct.
@@
where $n+1\leq i\leq m-3$. If $z_{i}$ and $z_{i-1}$ are equal, for example, $z_{i}=z_{i+2}=c$, then
@@
$f]=[k]=(t_{2}-t_{1})[e].$
Remark 3.9.
The computations for classical homotopy groups for posets rely heavily on the topological structures of spheres of arbitrary dimensions and their associated order complexes, and the continuous maps between them. It is known that to actually compute the classical homotopy groups is usually too difficult a procedure to be carried out in practice.
The process of obtaining the discrete homotopy groups of posets differs from the classical one in a significant way, as the last two examples show. The computation of these groups proceeds via the classification of all comparable maps in the category of posets, rather than via the homotopy classification of continuous maps in the category of topological spaces. This fact obviates the need for their associated order complexes. Problems regarding homotopy invariants of posets can now be tackled in a more visual and direct way in their own context.
@@ 4.1. Discrete cubical homology
Let $X$ be a finite poset. Denote by $Q_{1}$ the finite poset with vertex set $\{0,1\}$ and an edge $(0,1)$. For each $n\geq 1$, let $Q_{n}$ be the finite poset given by
@@
An $n$-cube of $X$ is a map $\sigma:Q_{n}\rightarrow X$.
Let $L_{n}(X)$ be the free abelian group generated by all $n$-cubes. We define two face maps from $L_{n}(X)$ to $L_{n-1}(X)$. Let $\sigma\in L_{n}(X)$. For $n\geq 1$, $i\in\{1,\cdots,n\}$ and $(a_{1},\cdots,a_{n-1})\in Q_{n-1}$, let $f_{i}^{-}\sigma$ and $f_{i}^{+}\sigma$ be the (n1)(n-1)-cubes of $X$ given by
@@
An $n$-cube $\sigma$ is degenerate if $f_{i}^{-}=f_{i}^{+}$ for some $i$. Let $D_{n}(X)$ be the free abelian group generated by all degenerate $n$-cubes and $C_{n}(X)=L_{n}(X)/D_{n}(X)$, whose elements are called $n$-chains in $X$.
@@
The boundary of an $n$-cube $\sigma:Q_{n}\rightarrow X$ is
@@
The boundary operator extends to a group homomorphism $\partial_{n}^{\rm Cube}:L_{n}(X)\rightarrow L_{n-1}(X)$. Since $\partial_{n}^{\rm Cube}(D_{n}(X))\subseteq D_{n-1}(X)$, we obtain a map $\partial_{n}^{\rm Cube}:C_{n}(X)\rightarrow C_{n-1}(X)$ and a chain complex $(C_{*}(X),\partial_{*}^{\rm Cube})$. Let
$H^{\rm Cube}_{n}(X)=\rm{Ker}\partial_{n}^{\rm Cube}/\rm{Im}\partial_{n+1}^{\rm Cube}.$
The group $H_{n}^{\rm Cube}(X)$ is called the $n$-th discrete cubical homology group of $X$. If $z\in\rm{Ker}\partial_{n}^{\rm Cube}$, then we write $[z]$ for the equivalence class of $z$ in $H_{n}^{\rm Cube}(X)$.
The connection between the discrete homotopy theory and the discrete homology theory is expressed by the following theorem.
@@
Let $X_{1}$ and $X_{2}$ be finite posets. If $X_{1}$ and $X_{2}$ are homotopy equivalent, then
$H^{\rm Cube}_{n}(X_{1})=H^{\rm Cube}_{n}(X_{2})$
@@
The idea of the proof of this theorem is exacly the same as the one in [4, Theorem 3.4]. We will not prove this theorem here.
@@
Let $K$ be a simplicial complex. For $n\geq 0$, we denote by $L^{\rm Simpl}_{n}(K)$ the free abelian group generated by all nondegenerate simplices of $K$ and write $D^{\rm Simpl}_{n}(K)$ for the free abelian group generated by all degenerate simplices of $K$. Let
$C^{\rm Simpl}_{n}(K)=L^{\rm Simpl}_{n}(K)/D^{\rm Simpl}_{n}(K).$
The $n$-th simplicial homology group of $K$ is denoted $H_{n}^{\rm Simpl}(K)$ whose associated chain complex is $(C^{\rm Simpl}_{*}(K),\partial_{*}^{\rm Simpl})$. Similarly, $H_{*}^{\rm Sing}$ denotes the singular homology theory. Denote by $\eta_{*}$ the canonical homomorphism
@@
To establish a natural homomorphism between the homology groups ${H}_{*}^{\rm Cube}(X)$ and ${H}_{*}^{\rm Simpl}(\mathcal{K}(X))$, we are motivated by the construction of the one in [4], which is a natural map from the cubical homology groups to the path homology groups. This map will be defined in a similar fashion.
Consider an $n$-cube $\sigma:Q_{n}\rightarrow X$ with $n\geq 1$. In order to define a map from ${C}_{n}^{\rm Cube}(X)$ to ${C}^{\rm Simpl}_{n}(\mathcal{K}(X))$, we associate to any permutation $\tau\in S_{n}$ an $n$-chain $p_{\tau}=\{x_{0},x_{1},\cdots,x_{n}~|~x_{0}<x_{1}<\cdots<x_{n}\}$ from $(0,\cdots,0)\in V(Q_{n})$ to $(1,\cdots,1)\in V(Q_{n})$. The chain $p_{\tau}$ is defined as follows. Let $x_{0}=(0,\cdots,0)$. For $0<i\leq n$, flipping the $\tau(i)$-th coordinate of $x_{i-1}$ from 0 to $1$ gives $x_{i}$. We write $p_{\tau}(i)$ for the $i$-th vertex in the chain $p_{\tau}$ for $0\leq i\leq n$.
To each $n$-cube $\sigma\in{L}_{n}^{\rm Cube}(X)$ for $n\geq 1$, we assign the chain
@@
of ${L}^{\rm Simpl}_{n}(\mathcal{K}(X))$. It is not hard to see that if $\sigma\in{D}_{n}^{\rm Cube}(X)$, then $\psi(\sigma)\in{D}^{\rm Simpl}_{n}(\mathcal{K}(X))$.
The proof that $\psi$ is a chain map is similar to the one in [4, Lemma 5.1]. For completeness, we will sketch this proof.
@@
Let $\sigma\in{C}_{n}^{\rm Cube}(X)$. Then
@@
$\psi_{*}:{H}_{p}^{\rm Cube}(X)\rightarrow{H}_{p}^{\rm Simpl}(\mathcal{K}(X))$
@@
Let $1\leq l\leq n-1.$ If $\tau^{\prime}$ is constructed from $\tau$ by interchanging $\tau(l)$ and $\tau(l+1)$, then $p_{\tau}(i)=p_{\tau^{\prime}}(i)$ for $i\not=l$. Thus the $l$-th summands of (4.1) for $p_{\tau}$ and $p_{\tau^{\prime}}$ coincide. In addition, ${\rm sign}(\tau)={\rm-sign}(\tau^{\prime})$. This means that
@@
Let $X$ be a homogeneous poset of dimension $1$. The map $\psi$ induces an isomorphism $\psi_{*}:{H}_{p}^{\rm Cube}(X)\rightarrow{H}_{p}^{\rm Simpl}(\mathcal{K}(X))$ for any $p\geq 0$.